OpenStudy (anonymous):

Solve for x 5x-2+3x=5x-9

hartnn (hartnn):

collect like terms, terms with all x's together

OpenStudy (anonymous):

8x-5x=-9+2 3x=-7 x=-7/3 or -2 1/3.

hartnn (hartnn):

for that u need to bring 5x of right to left side, u know how to ?

OpenStudy (anonymous):

so i get all the x's on one side of the equation first?

hartnn (hartnn):

yup , and all constants (numbers) on other side

OpenStudy (anonymous):

ok, so when i do that my equation looks like this :8x-5x=-9+2 Next do i subtract 5x from 8x?

hartnn (hartnn):

yes, whats 8x-5x = ?

OpenStudy (anonymous):

3x right ?

hartnn (hartnn):

yes : ) and whats right side -9+2= ?

OpenStudy (anonymous):

-7 ?

hartnn (hartnn):

so u get 3x=-7 what next ?

OpenStudy (anonymous):

Do i divide both sides by 3 to get x alone ?

hartnn (hartnn):

that is correct! you are good learner what u get after dividing ?

OpenStudy (anonymous):

Thanks, x=-2.33

OpenStudy (anonymous):

thats what i got.. is it right?

hartnn (hartnn):

yes, that is right, or you can keep it as -7/3 also :)

OpenStudy (anonymous):

ok thanks :)

hartnn (hartnn):

welcome ^_^